OpenStudy (anonymous):

There are two consecutive odd numbers such that twice the smaller number plus one is ten more than the greater number. What is the smaller number? 7 9 11 13

OpenStudy (ajprincess):

Let the larger number be x and smaller one be y. x=y+2 x+10=2y+1 x=2y+1-10 x=2y-9 2y-9=y+2 2y-y=9+2 y=11 Is that clear?

OpenStudy (anonymous):

thank you

OpenStudy (ajprincess):

yw